# RomanLens: Latent Romanization and its role in Multilinguality in LLMs

### v1

@@ 7 Conclusion
In our investigation of romanized representations in the latent layers, we conclusively identified romanized tokens in the last 6–7 layers of an LLM across various multilingual text completion tasks. Based on previous works in this field Wendler et al. (2024); Zhao et al. (2024b), in an English-centric decoder only LLM this region corresponds to the transition from an English-centric language-agnostic concept space to a language-specific space where the idea conceived in the concept space is expressed in the target language. Our findings suggest that romanization serves as a bridge between the concept space and the language-specific region for non-Latin script languages, an observation strongly supported by our analysis of six diverse writing systems, though notably not observed in Chinese. Romanization acting as a bridge could explain why romanization based script barrier breaking methods like Liu et al. (2024); Xhelili et al. (2024) work.
Latent romanization is not observed for Chinese, unlike the other six languages examined in this study. This difference may be partially attributed to Chinese being a relatively high-resource language compared to the others analyzed, as well as its logographic script.
The identification of latent romanization in LLMs underscores its potential as a practical mechanism to address the challenges of limited training data for under-resourced languages. This perspective shifts the narrative from viewing such phenomena as an inherent English-centric bias to recognizing them as opportunities for innovation in multilingual AI. While our study provides initial insights into latent romanization, further research would be needed to understand whether and how these findings could inform training strategies for improving language model performance across diverse linguistic communities.
@@ 8 Limitations
The handling of multilingual text by large language models (LLMs) is an ongoing area of research. While evidence suggests that LLMs interact with English representations in a language-agnostic space, the mechanisms governing how this interaction varies at different timesteps during token generation of a word remain unclear. Our findings indicate that romanized representations become more prominent in hidden layers as the model progresses from the first to the last token generation step. This suggests that romanization may help the model address token fertility differences between output language and its primary latent language, English, especially for non-Latin languages with high token fertility. However, further research is needed to confirm and generalize these observations.
Furthermore, the scope of our analysis is restricted to models that utilize SentencePiece based tokenizers. The interpretability of non-Latin scripts at latent layers is limited when models employ tokenization schemes that split characters into multiple bytes, complicating logit lens analysis. Extending this work to models with alternative tokenization methods would offer a more complete understanding of multilingual capabilities and representations.
@@ Acknowledgments
We would like to thank EkStep Foundation and Nilekani Philanthropies for their generous grant towards research at AI4Bharat.
Appendix A English Translation and Romanization of the Sample Prompts
English translation and romanizations for the sample translation and repetition prompts mentioned in 4.1 are provided below.
@@ Appendix A English Translation and Romanization of the Sample Prompts
Appendix B Latent romanization: Tokenization scheme for the romanized word
@@
Appendix C Latent Fraction
@@
Appendix D Computing Language probabilities - For translation towards native script vs translation towards romanized script task
@@
Appendix E Computing Probabilities : Activation Patching Experiment
@@
Appendix F Latent Romanization Qualitative Analysis
We list qualitative logit lens analysis for Greek, Ukrainian, Hebrew, Arabic and Amharic (see Figures 9 to 13).
@@ Appendix F Latent Romanization Qualitative Analysis
Appendix G Latent Romanization Quantitative Analysis
Quantitative Analysis of latent romanization for repetition task with gemma 2 9b it model can be seen in Figure 14. Layerwise fractional distribution of romanized tokens across output token generation timesteps for translation and repetition task with Gemma 2 9b, Llama 2 7b, and Llama 2 13b models are present in Figure 15.
@@ Appendix G Latent Romanization Quantitative Analysis
Appendix H Translation towards native script vs Translation towards romanized script: Additional examples

### v2

@@ 7 Conclusion
Our findings show that LLMs implicitly use Romanization as a bridge for non-Roman scripts, exhibiting Latent Romanization in intermediate layers before switching to native scripts. Layerwise analyses reveal that semantic concepts are encoded similarly across native and Romanized inputs, indicating a shared internal representation. Moreover, when translating to a Romanized script, target words emerge earlier, highlighting Romanization as a structural link between language-agnostic concepts and language-specific output. While our study reveals initial insights into Latent Romanization, future work could focus on applying these findings to develop training strategies that enhance performance across diverse linguistic communities.
@@ 8 Limitations
The handling of multilingual text by large language models (LLMs) remains an active area of research. Although evidence suggests that LLMs process English representations within a language-agnostic space, the specific mechanisms by which these models adjust their interactions over different timesteps during token generation are still not fully understood. In our study, we observe that romanized representations become increasingly prominent in the hidden layers as token generation progresses from the first to the final token. This trend suggests that latent romanization may help the model mitigate differences in token fertility—that is, the average number of tokens required to represent a word—between the output language and its primary latent language, English. This effect appears especially for non-Roman script languages, with high token fertility. However, further research is needed to confirm and generalize these observations.
The interpretability of non-Roman scripts at latent layers is limited when models employ tokenization schemes that split non-roman characters into multiple bytes, complicating logit lens analysis. Extending this work to models with alternative tokenization methods would offer a more complete understanding of multilingual capabilities and representations.
@@ 9 Ethics Statement
The code and datasets created in this work will be made available under permissible licenses. Generative AI systems were only used for assistance purely with the language of the paper, eg: paraphrasing, spell-check, polishing the author’s original content, and for writing boiler-plate code.
@@ Acknowledgments
Appendix A Transformer’s Forward pass: Detailed
For an input sequence , where is the sequence length, the initial latents are obtained from a learned embedding matrix. The update rule for the latent at position in layer is expressed as:
The logit scores are computed as:
These are converted to probabilities via the softmax function:
Appendix B Sample Prompts
A Hindi example, its English translation and transliteration for the translation, repetition and cloze task prompt designs mentioned in Section 4 are provided below.
Hindi example.
Translation task. A translation prompt from French to Hindi.
[…]
Repetition task.
[…]
Cloze task..
[…]
@@ Appendix B Sample Prompts
Appendix C Latent romanization: Tokenization scheme for the romanized word
@@
Appendix D Latent Fraction
@@
Appendix E Computing Language probabilities - For translation towards native script vs translation towards romanized script task
@@
Appendix F Romanization scheme: Indic languages
We have taken into consideration two romanization schemes for Indic languages: (a) ITRANS scheme from IndicNLP library Kunchukuttan (2020) and (b) IndicXlit scheme Madhani et al. (2023). Based on our initial experiments, we observed that the IndicXlit scheme produces better romanization than ITRANS scheme. Thus for romanization we have used the IndicXlit scheme Madhani et al. (2023). It generates romanization as is commonly used by native speakers and is trained on parallel transliteration corpora.
Appendix G Computing Probabilities : Activation Patching Experiment
@@
Appendix H Latent Romanization Qualitative Analysis
We list qualitative logit lens analysis for Greek, Ukrainian, Hebrew, Arabic and Amharic (see Figures 8 to 12).
Languages.
Greek is part of the Hellenic branch of the Indo-European language family and is written using the Greek alphabet. Ukrainian belongs to the East Slavic group of the Indo-European family and employs the Ukrainian alphabet, a variant of the Cyrillic script. Amharic is a South Ethio-Semitic language within the Afroasiatic family and is written using the Ge’ez script, an abugida where each character represents a consonant-vowel combination. Hebrew is a Northwest Semitic language within the Afroasiatic family and is written using the Hebrew alphabet, an abjad script originating from the Aramaic alphabet. Arabic is a Central Semitic language, also part of the Afroasiatic family, and utilizes the Arabic script, another abjad that evolved from the Nabataean alphabet. Notably, both Hebrew and Arabic scripts are written from right to left.
@@ Appendix H Latent Romanization Qualitative Analysis
Appendix I Latent Romanization Quantitative Analysis: Additional examples
Quantitative Analysis of latent romanization for repetition task and cloze task with gemma 2 9b it model can be seen in Figures 13 and 14 respectively. Layerwise fractional distribution of romanized tokens across output token generation timesteps for translation, repetition and cloze task with Gemma 2 9b, Llama 2 7b, and Llama 2 13b models are present in Figure 15.
@@ Appendix I Latent Romanization Quantitative Analysis: Additional examples
Cloze
Appendix J Comparing Translations Into Romanized vs. Native Script: Additional examples
@@
Appendix K Other Models: Mistral
We also perform our experiments on Mistral-7B Jiang et al. (2023), a popular LLM known for its performance and efficiency. Layerrwise distribution of romanized tokens for initial, final and all token generation steps are presented in Figure 19.
| All tokens | First token | Last token |
Translation
Repetition
Cloze